Use the fact [x y 1]

[a1 b1 1] =0

[a2 b2 1]

To determine the equation of the line passing through the distinct points(a1, b1) and (a2, b2) where (.) stands for det.


1
Expert's answer
2021-06-07T13:27:24-0400


We have,

"\begin{vmatrix}
 x & x&1 \\
 a_1 & b_1&1\\
 a_2&b_2&1
\end{vmatrix}=0\\\ \\\Rightarrow x(b_1-b_2)-y(a_1-a_2)+b_2a_1-a_2b_1=0\\\Rightarrow x(b_2-b_1)-y(a_2-a_1)+a_2b_1-b_2a_1=0\\\Rightarrow x(b_2-b_1)+a_2b_1=y(a_2-a_1)+b_2a_1\\\text{Add }a_2b_2\ on\ both\ sides,\ we\ get\\\Rightarrow x(b_2-b_1)-a_2b_2+a_2b_1=y(a_2-a_1)-b_2(a_2-a_1)\\\Rightarrow (b_2-b_1)(x-a_2)=(a_2-a_1)(y-b_2)\\ \ \\ \Rightarrow\dfrac{y-b_2}{b_2-b_1}=\dfrac{x-a_2}{a_2-a_1}"